A HClO / NaClO buffer system is to be constructed that is also 0.042 M in Co2+. What is the minimum [HClO] / [ClO-] that must be maintained to keep a precipitate of cobalt(II) hydroxide from forming?

minimum [HClO] / [ClO-]= ___?___ / 1
Answer #1

Sol . As For Cobalt (II) hydroxide ( Co(OH)2 ) ,

Co(OH)2 <----> Co2+ + 2OH-

So , Ksp = [Co2+] [OH-]2

As Ksp = 3.3 × 10-16 and [Co2+] = 0.042 M

So , 3.3 × 10-16 = 0.042 × [OH]2

and , [OH-] = ( 78.5714 × 10-16 )1/2 = 8.8640 × 10-8 M

Also , pOH = - log[OH-] = - log ( 8.8640 × 10-8 ) = 7.05

Therefore , pH of buffer = 14 - pOH = 14 - 7.05 = 6.95  

Now , pKa of HClO/NaClO buffer = 7.46

So , Using Henderson - Hasselbalch equation ,

pH = pKa + log ( [NaClO] / [HClO] )  

6.95 = 7.46 + log ( [NaClO] / [HClO] )

log ( [NaClO] / [HClO] ) = - 0.51

So , [NaClO] / [HClO] = 10-0.51 = 0.3090

and , [HClO] / [NaClO] = [HClO] / [ClO-] = 1 / 0.3090 =  3.2362